# Gravitational-Wave Signatures of Highly Eccentric Stellar-Mass Binary Black Holes in Galactic Nuclei

### v1

@@ 3.4 Merging orbits
Fig. 6 shows the orbital evolution of the merging orbit, which is the most varied of the orbital evolution families and has several phases. Initially, the orbit is in libration. However, due to the initial large eccentricity, each pericentre passage increases the libration amplitude: the energy decay changes the locations of the fixed point and detunes the system, making the libration amplitude larger. The decreasing semi-major axis causes the secular period to increase, and eventually, after $\sim 6$ cycles and about $\sim 30$ years of orbital evolution, $\omega$ escapes and starts to circulate. After another cycle, GW capture occurs and the GW dissipation is strong enough to reduce the semi-major axis such that the inner binary is essentially decoupled from the central BH.
Finally, the inner binary merges. Using Peters (1964) to evolve the binary from its final recorded state at $0.26$ Hz, we estimate that the eccentricity of this binary at $10$ Hz 22-mode GW frequency is $0.016$. This may be slightly below the current threshold of LVK detectors ($0.05$ at $10$ Hz for GW150914-like systems detected with third observing run sensitivity; Lower et al., 2018), however, lighter systems detected with increasingly sensitive detectors may have a lower minimum detectable eccentricity. Future third-generation detectors with improved sensitivity and lower minimum frequencies, expected in the mid-2030s and coinciding with LISA, will improve eccentricity measurement sensitivity significantly (Saini, 2024). At $1$ Hz ($5$ Hz) 22-mode GW frequency this binary retains an eccentricity of $0.172$ ($0.033$). The rapid variations of $\omega$ observed in Fig. 6 are due to fast GR precession from the 11PN term, which is associated with the increase of the GW frequency, and potentially observable in ground-based GW detectors. We also note that the final inclination with respect to the outer orbit is retrograde, exceeding $100$ deg.
Figure 7 shows the evolution of the merging orbit in the $a-e_{1}$ parameter space (top panels) and and in $f_{\rm GW}-e_{1}$ space (bottom panels), colour coded by the time before merger $t_{\rm merge}-t$. We see that initially the binary is at fixed $a$ and emits a burst of GW radiation every time it becomes highly eccentric, near $1-e\sim 10^{-3}\times(0.6-2)$, causing it to experience a step-like decrease in separation. After about $\mathcal{O}(10)$ cycles, the binary is eventually decouples from the tertiary SMBH.
@@
The two effects start to dominate over the ZLK oscillations roughly on the same time. Using the Peters (1964) formula for the GW timescale at the 2.5PN level and the timescale for the 1PN precession, the timescale ratio is
@@ 4.1 Distinguishing between the orbits
The merging orbit is the most probable observable GW source. This orbit starts on a low value of $j_{z}$, which leads to a high eccentricity and corresponding $f_{\rm GW}$. In the first few cycles of the GW capture, $\omega$ rapidly circulates due to the 1PN term; this occurs at $f_{\rm GW}\approx 0.3\ \rm Hz$. The simulation stops at $f_{\rm GW}\approx 0.48\ \rm Hz$ at $a_{0}$ and $e_{0}$, as in Fig. 7. The merging orbit may be observed in LISA first as a high-eccentricity burst source, before being captured inspiralling through LISA and into ground-based GW detectors. This source may be observed in real-time in LISA, or alternatively may be observed with ground-based detectors before being searched for in archival LISA data.
We compare the different strains obtained for our four representative orbits in Figure 8. We use the same sampling rate of $0.0001$ Hz and the same time of integration of $10$ years. We see that for the orbits that exhibit large eccentricity variation—i.e, all orbital families except small libration—the strain has distinct sharp bursts and reaches strain amplitude magnitudes up to $\approx 10^{-19}$, and higher for the merging binary (for the Galactic distance of $8\ \rm kpc$). The time difference between the peaks is the secular time, which is slightly different between the orbits. The merging orbit has a varying secular timescale due to the changing libration amplitude (the capture phase is not shown). The small librating orbit does not exhibit strong bursts and is more sinusoidal throughout the evolution.
These waveforms are generated numerically using the new functionality we have added to the TSUNAMI code. If the time between output snapshots is $\Delta t$, than each waveform takes $\mathcal{O}(1)\times(1+\Delta t/5000\ \rm s)$ s to generate where, which makes Bayesian inference studies infeasible. In future work, we will extract burst time information from TSUNAMI and create an approximate surrogate model to enable the parameters of the systems to be inferred (e.g., Romero-Shaw et al., 2023a).
4.2 Generalization to arbitrary tertiary mass
@@
Figure 9 shows the evolution of the small libration orbit for different values of $\gamma$. We see that the equal mass case ($\gamma=0.0171;\ m_{\rm out}=\gamma^{3}\times 4\times 10^{6}\approx 20\ M_{\odot}$) is evolving significantly differently, as seen in the orbital elements in the top panels. The difference between the burst timing (bottom middle panel) is about $3$ hours. The other orbits ($\gamma=0.8,0.99$) evolve similarly to the fiducial orbit, and the orbital element evolution is indistinguishable after 1 year. However, since the width of each burst is narrow, the GW strain signature is still different. The bottom right panel shows that the offset between the $\gamma=0.8$ orbit and the fiducial orbit is about $\approx 4-5\ \rm mins$, while the typical width of the burst is about $2\ \rm mins$, so the burst are largely separate from each other. Only the $\gamma=0.99$ orbit significantly overlaps with the fiducial orbit, with an offset of about $\sim 20\ \rm sec$. However, one year of data contains over $100$ peaks (since the inner orbital period is slightly less than $0.01\ \rm yr$). Even orbits with small differences in $\gamma$ are therefore expected to be distinguishable.
@@ 4.3 Comparison to an isolated binary
Figure 10 shows the evolution of the small libration orbit compared to an isolated eccentric binary (without the SMBH). We test two initial values of the eccentricity: the same as in the small libration case (at the fixed point; dashed blue) and at a lower value closer to the mean around the secular oscillations (dashed green). We see that in the isolated cases $\omega$ precesses due to the 1PN term, and a larger eccentricity leads to faster precession, due to the $(1-e_{1}^{2})^{-}1$ term in the 1PN precession rate. The eccentricity in the isolated case is almost fixed due to lack of perturbations, contrary to the small oscillations in the SMBH case. Similar behaviour is evident in $f_{\rm GW}$ (bottom panels) where the GW frequency is varying mostly due to the eccentricity oscillations but remains largely fixed for the isolated binaries. The strain polarirsaion looks dense with characteristic timescale of modulation either with the ZLK oscillations (SMBH case) or due to the 1PN precession (isolated cases). We also note that indeed the secular timescale is much faster than the 1PN precession, which makes $\epsilon_{\rm GR}$ (Eq. 2) small, as expected.
Besides the 1PN precession timescale and secular ZLK timescale, we also observe additional timescale in the time series. First, the strain polarisation is bursty with typical time between burst being the inne period of $P_{1}\approx 0.00918\ \rm yr=3.36\ \rm days$. Small ’sawtooth’ oscillations in the eccentricity and $f_{\rm GW}$ are evident on this timescale. The eccentricity also oscillates on a longer timescale of $P_{2}\approx 0.23\ \rm yr$. At every pericentre, the orbital elements change rapidly and the strain increases by several orders of magnitude for a short duration of $\sim r_{p}/v_{p}$ due to the large accelerations involved. Zooming in on one peak, we see that the arrival times (and also strain amplitudes) are well separated by about 2 hours from each other, which is about $\sim 2.5\%$ of the total orbital period. Thus, the different orbits can be well separated from each other.
@@ 5 Summary and Conclusions
We have explored the complex behaviour of BBHs in galactic nuclei influenced primary by the tertiary SMBH, in the process adding to the TSUNAMI code the capability to produce strain signals. Our main conclusions are as follows:
@@
$\bullet$ In addition to the standard von Zeipel-Lidov-Kozai (ZLK) evolution where the inner argument of pericentre $\omega$ is circulating (Fig. 3), $\omega$ and the eccentricity $e_{1}$ can also librate with large (Fig. 4) or small (Fig. 5) amplitude around a fixed point in phase space, which is given by Eq. (3) and $\omega=\pi/2,3\pi/2$ (see also Grishin, 2024a). Moreover, for extremely high eccentricity (or low $j_{z}$), GW emission at pericentre can lead to GW capture and merger even an initially librating orbit (Fig. 6), which could potentially be observed both by LISA and ground-based GW detectors. These four families have unique signatures in $\omega$ and GW strain (Fig. 2).
@@
$\bullet$ We added a method to calculate the strain polarisations within the N-body code TSUNAMI (Trani & Spera, 2023) by using the quadrupole formula and accounting for the direct accelerations from the tertiary and due to post-Newtonian effects up to 3.5PN. Using this method we are able to compare the different GW signatures directly from the simulations (Fig. 8). We find that most of the strain signatures include high bursts separated by a quiescent phase on timescales of years. The small-amplitude orbit maintains its high eccentricity and has consistent burst on the order of its inner orbital period of a few days.
@@
$\bullet$ Due to the complexity of the dynamics, each orbit has almost unique signature. Varying the SMBH mass by as little as $3\%$ ($\gamma=0.99$ in Fig. 9) will lead to an offset of the time-of-arrival of the burst by avout $20$ secs, which can still enable distinguishability, based on the precision to which we are able to isolate the arrival time of a single burst. Distinguishing BBH in galactic nuclei from those formed in more equal-mass field triples is also possible due to the breakdown of the test-particle limit and the backreaction on the tertiary orbit (Fig. 9). Moreover, the orbits of BBHs with a SMBH perturber are markedly different from a binary in isolation with the same initial conditions (Fig. 10).
The rates of the systems we investigate in this paper are uncertain. Although about $10^{4}$ BHs are expected to reside within a distance of $10^{3}\ \rm au$ from Sgr A* (Miralda-Escudé & Gould, 2000; Freitag et al., 2006; Rose et al., 2022), it is unclear how many of them are in binaries or in librating configuration. Taking an example from the Solar system, of the $228$ irregular satellites around giant planets, only only $13/228\approx 5.7\%$ are found to be librating (Grishin, 2024b). BBHs, of course, have very different formation and evolution channels. If the BBH binary fraction is $f_{\rm bin}$, and the mutual inclination is uniform in cosine, we may expect $\sim 10^{4}f_{\rm bin}f_{\rm high}$, where $f_{\rm high}=0.1$ is the number of binaries with $|j_{z}|\leq 0.1$. Taking the fiducial $5.7\%$ value of librating orbits, we end up with $\sim 57f_{\rm bin}$ librating orbits with high inclination (and eccentricity). The binary fraction is unknown, but BHs do tend to pair up efficiently in nuclear star clusters (e.g., Kritos et al., 2023; Chattopadhyay et al., 2023; Dall’Amico et al., 2024).
@@
The duration of a burst is about $r_{p}/v_{p}$ at periapse, which is $\approx 1$ minute, as in Fig. 9. This occurs every $\sim 3.4$ days, or >4800>4800 minutes. Generating 1 year of data takes $\approx 0.5$ seconds for output snapshots above $\Delta t=5,000$ sec, and linearly increases to $\sim 4$ minutes for the desired output of $\Delta t=10\ \rm s$ in order to capture the burst time derivative. Thus, full Bayesian parameter estimation is not feasible using the waveform model we have developed. However, searches and inference methods that rely on bursts and burst timing, rather than generation of full waveforms, can identify these sources and potentially infer the properties of the system (Romero-Shaw et al., 2023a; Knee et al., 2024; Saini et al., 2025).

### v2

@@ 3.4 Merging orbits
Figure 6 shows the orbital evolution of the merging orbit, which is the most varied of the orbital evolution families and has several phases. Initially, the orbit is in libration. However, due to the initial large eccentricity, each pericentre passage increases the libration amplitude: the energy decay changes the locations of the fixed point and detunes the system, making the libration amplitude larger. The decreasing semi-major axis causes the secular period to increase, and eventually, after $\sim 6$ cycles and about $\sim 30$ years of orbital evolution, $\omega$ escapes and starts to circulate. After another cycle, GW capture occurs and the GW dissipation is strong enough to reduce the semi-major axis such that the inner binary is essentially decoupled from the central BH.
Finally, the inner binary merges. Using Peters (1964) to evolve the binary from its final recorded state at $0.26$ Hz, we estimate that the eccentricity of this binary at $10$ Hz 22-mode GW frequency is $0.016$. This may be slightly below the current threshold of LVK detectors ($0.05$ at $10$ Hz for GW150914-like systems detected with third observing run sensitivity; Lower et al., 2018), however, lighter systems detected with increasingly sensitive detectors may have a lower minimum detectable eccentricity. Future third-generation detectors with improved sensitivity and lower minimum frequencies, expected in the mid-2030s and coinciding with LISA, will improve eccentricity measurement sensitivity significantly (Saini, 2024). At $1$ Hz ($5$ Hz) 22-mode GW frequency this binary retains an eccentricity of $0.172$ ($0.033$). The rapid variations of $\omega$ observed in Figure 6 are due to fast GR precession from the 11PN term, which is associated with the increase of the GW frequency, and potentially observable in ground-based GW detectors. We also note that the final inclination with respect to the outer orbit is retrograde, exceeding $100$ deg.
Figure 7 shows the evolution of the merging orbit in the $a-e_{1}$ parameter space (top panels) and in $f_{\rm GW}-e_{1}$ space (bottom panels), colour coded by the time before merger $t_{\rm merge}-t$. We see that initially the binary is at fixed $a$ and emits a burst of GW radiation every time it becomes highly eccentric, near $1-e\sim 10^{-3}\times(0.6-2)$, causing it to experience a step-like decrease in separation. After about $\mathcal{O}(10)$ cycles, the binary eventually decouples from the tertiary SMBH.
@@
The two effects start to dominate over the ZLK oscillations roughly at the same time. Using the Peters (1964) formula for the GW timescale at the 2.5PN level and the timescale for the 1PN precession, the timescale ratio is
@@ 4.1 Distinguishing between the orbits
Figure 9: Characteristic strain $h_{c}=2f|\tilde{h}(f)|$ of the four orbital families overlaid with the LISA’s sensitivity curve (thick black line). The signals are here curtailed artificially suddenly at the highest-frequency point in the simulation output.
The merging orbit is the most probable observable GW source. This orbit starts on a low value of $j_{z}$, which leads to a high eccentricity and corresponding $f_{\rm GW}$. In the first few cycles of the GW capture, $\omega$ rapidly circulates due to the 1PN term; this occurs at $f_{\rm GW}\approx 0.3\ \rm Hz$. The simulation stops at $f_{\rm GW}\approx 0.48\ \rm Hz$ at $a_{0}$ and $e_{0}$, as in Figure 7. The merging orbit may be observed in LISA first as a high-eccentricity burst source, before being captured inspiraling through LISA and into ground-based GW detectors. This source may be observed in real-time in LISA, or alternatively may be observed with ground-based detectors before being searched for in archival LISA data.
We compare the different strains obtained for our four representative orbits in Figure 8. We use the same sampling rate of $10^{-4}$ Hz and the same time of integration of $10$ years. We see that for the orbits that exhibit large eccentricity variation—i.e, all orbital families except small libration—the strain has distinct sharp bursts and reaches strain amplitude magnitudes up to $\approx 10^{-19}$, and higher for the merging binary (for the Galactic distance of $8\ \rm kpc$). The time difference between the peaks is the secular time, which is slightly different between the orbits. The merging orbit has a varying secular timescale due to the changing libration amplitude (the capture phase is not shown). The small librating orbit does not exhibit strong bursts and is more sinusoidal throughout the evolution.
Figure 9 shows the characteristic strain of the four families overlaid with the LISA sensitivity curve. The characteristic strain is $h_{c}(f)=2f|\tilde{h}(f)$|, where $\tilde{h}(f)=\sqrt{\tilde{h}^{2}_{x}(f)+\tilde{h}^{2}_{+}(f)}$ is the sum of the Fourier components of the two strain polarisations. For the Fourier transform we use the fast Fourier Transform implemented within the BILBY package (Ashton et al., 2019), where scipy.interpolate’s interpolator was used to construct a uniform time series from the simulated strains. LISA’s sensitivity curve (thick black line) was generated using Robson, Cornish, & Liu (2019)’s Jupyter notebook.
We see that all orbital families have a distinctive detectable features where the peak strain exceeds the LISA sensitivity curve by about two orders of magnitude near the minimum of the sensitivity curve around $f_{\rm GW}\approx 3-5\times 10^{-3}\ \rm Hz$. The merging orbit remains high throughout the LISA frequency band. These preliminary high SNRs make future detections potentially feasible.
These waveforms are generated numerically using the new functionality we have added to the TSUNAMI code. If the time between output snapshots is $\Delta t$, then each waveform takes $\mathcal{O}(1)\times(1+\Delta t/5000\ \rm s)$ s to generate, which makes Bayesian inference studies infeasible. In future work, we will extract burst time information from TSUNAMI and create an approximate surrogate model to enable the parameters of the systems to be inferred (e.g., Romero-Shaw et al., 2023a).
4.2 Total number of librating BBH
The rates of the systems we investigate in this paper are uncertain. Although about $10^{4}$ BHs are expected to reside within a distance of $10^{3}\ \rm au$ from Sgr A* (Miralda-Escudé & Gould, 2000; Freitag et al., 2006; Rose et al., 2022), it is unclear how many of them are in binaries or in librating configuration. The only large statistical sample of well resolved orbits of similar configurations of small-mass binaries orbits a much more massive object is the sample of irregular satellites in the Solar system, where $228$ irregular satellites around giant planets are known, and only $13/228\approx 5.7\%$ are found to be librating (Grishin, 2024b). However, these orbits are markedly different with relative low inclinations with $|j_{z}|>0.4$, so they do not experience high eccentricities. There is also a distinct void in the range of mutual inclinations $60-120^{\circ}$, attributed to the loss of the satellite during ZLK cycles (Grishin et al., 2017).
The conditions for libration to be possible is given by the dimensionless quantities $\epsilon_{\rm SA}$ (Eq. 3) and $j_{z}$ (see Figure 7 of Grishin, 2024a for ’optimal’ libration zone where the initial $\omega=\pi/2$ and Figure 1 of Grishin, 2024b for the sample of known irregular satellites), however they are inaccurate since they neglect the dependence of $\omega$.
Recently, a high-order analytical theory was developed to classify librating orbits with the extent of the Brown Hamiltonian (accounting for mild hierarchies, Lei & Grishin, 2025a, b). In particular, the orbit will be librating if the modified ’Kozai constant’
$C_{\rm ZLK}=e^{2}\left[1+\frac{9}{2}\epsilon_{\rm SA}j_{z}-\frac{5}{2}\left(1+\frac{9}{8}\epsilon_{\rm SA}j_{z}\right)\sin^{2}\omega\sin^{2}\iota\right]$ (12)
is negative (Eq. 20 in Lei & Grishin, 2025b, see also their Figure 2 for parameter space exploration). For $\epsilon_{\rm SA}=0$, $C_{\rm ZLK}$ reduces to the classical ZLK constant.
Figure 10: Parameter space of $C_{\rm ZLK}$-$j_{z}$ that distinguishes circulating ($C_{\rm ZLK}>0$) and librating $(C_{\rm ZLK}<0)$ orbits from our fiducial sample. About $37\%$ of the orbits are librating (top panel CDF). Another $\sim 20\%$ has low $j_{z}$ (bottom panel CDF).
@@ 4.2 Total number of librating BBH
BBHs, of course, have very different formation and evolution channels. If the BBH binary fraction is $f_{\rm bin}$, and the mutual inclination is uniform in cosine, we may expect $\sim 10^{4}f_{\rm bin}f_{\rm high,lib}$ where $f_{\rm high,lib}$ is the number of librating binaries with $|j_{z}|\leq 0.1$ such that hte eccentricity is large.
In order to explore these fractions, we take $N=10^{6}$ random samples of a triple system with $\omega\sim U[0,2\pi]$ sampled from a uniform distribution, $e$ sampled from a thermal distribution $(f(e)=2e).\cos\iota\sim U[-1,1]$ also uniform and $\log_{\rm 10}\epsilon_{\rm SA}\sim U[-3,-0.5]$ is sampled from a log-uniform distribution corresponding to Öpik’s law.
Figure 10 shows the distribution of the orbits in $j_{z}$-$C_{\rm ZLK}$ space, which is very similar to the parameter space presented in Figure 2 of Lei & Grishin (2025b). We see that $37\%$ of the orbits are librating $(C_{\rm ZLK}<0)$, $\sim 20\%$ have $|j_{z}|>0.1$. The total number of highly-eccentric librating orbits (both $C_{\rm ZLK}<0$ and $|j_{z}|<0.1$) is about $10\%$. For the irregular satellite population, about $7\%$ of the orbits are librating with $|j_{z}|>0.4$, compatible with the observations. Fixing $\epsilon_{\rm SA}=10^{-0.5}$ only raises the fraction of librating orbit to $\sim 42\%$ but the fraction of highly-eccentric librating orbits remains similar, thus we conclude that our studies sample does not strongly depend on the prior of $\epsilon_{\rm SA}$.
To summarise, $f_{\rm high,lib}\approx 0.1$ and we expect about $10^{3}f_{\rm bin}$ highly eccentric librating binaries. The same estimate for low eccentricity librating binaries is about $7\%$, compatible with the $5.7\%$ of the librating irregular satellite sample. The binary fraction is unknown, but BHs do tend to pair up efficiently in nuclear star clusters (e.g., Kritos et al., 2023; Chattopadhyay et al., 2023; Dall’Amico et al., 2024).
4.3 Generalisation to arbitrary tertiary mass
@@
Figure 11 shows the evolution of the small libration orbit for different values of $\gamma$. We see that the equal mass case ($\gamma=0.0171;\ m_{\rm out}=\gamma^{3}\times 4\times 10^{6}\approx 20\ M_{\odot}$) is evolving significantly differently, as seen in the orbital elements in the top panels. The difference between the burst timing (bottom middle panel) is about $3$ hours. The other orbits ($\gamma=0.8,0.99$) evolve similarly to the fiducial orbit, and the orbital element evolution is indistinguishable after 1 year. However, since the width of each burst is narrow, the GW strain signature is still different. The bottom right panel shows that the offset between the $\gamma=0.8$ orbit and the fiducial orbit is about $4-5\ \rm min$, while the typical width of the burst is about $2\ \rm mins$, so the bursts are largely separate from each other. Only the $\gamma=0.99$ orbit significantly overlaps with the fiducial orbit, with an offset of about $20\ \rm s$. However, one year of data contains over $100$ peaks (since the inner orbital period is slightly less than $0.01\ \rm yr$). Even orbits with small differences in $\gamma$ are therefore expected to be distinguishable.
@@ 4.4 Comparison to an isolated binary
Figure 12 shows the evolution of the small libration orbit compared to an isolated eccentric binary (without the SMBH). We also use the isolated case as a benchmark for the strain calculation which show a smooth modulation on top of the burst at each pericentre passage. We test two initial values of the eccentricity: the same as in the small libration case (at the fixed point; dashed blue) and at a lower value closer to the mean around the secular oscillations (dashed green). We see that in the isolated cases $\omega$ precesses due to the 1PN term, and a larger eccentricity leads to faster precession, due to the $(1-e_{1}^{2})^{-1}$ term in the 1PN precession rate. The eccentricity in the isolated case is almost fixed due to lack of perturbations, contrary to the small oscillations in the SMBH case. Similar behaviour is evident in $f_{\rm GW}$ (bottom panels) where the GW frequency is varying mostly due to the eccentricity oscillations but remains largely fixed for the isolated binaries. The strain polarisation looks dense with characteristic timescale of modulation either with the ZLK oscillations (SMBH case) or due to the 1PN precession (isolated cases). We also note that indeed the secular timescale is much faster than the 1PN precession, which makes $\epsilon_{\rm GR}$ (Eq. 2) small, as expected.
Besides the 1PN precession timescale and secular ZLK timescale, we also observe additional timescale in the time series. First, the strain polarisation is bursty with typical time between burst being the inner period of $P_{1}\approx 0.00918\ \rm yr=3.36\ \rm days$. Small ’sawtooth’ oscillations in the eccentricity and $f_{\rm GW}$ are evident on this timescale. The eccentricity also oscillates on a longer timescale of $P_{2}\approx 0.23\ \rm yr$. At every pericentre, the orbital elements change rapidly and the strain increases by several orders of magnitude for a short duration of $\sim r_{p}/v_{p}$ due to the large accelerations involved. Zooming in on one peak, we see that the arrival times (and also strain amplitudes) are well separated by about 2 hours from each other, which is about $\sim 2.5\%$ of the total orbital period. Thus, the different orbits can be well separated from each other.
@@ 5 Summary and Conclusions
We have explored the complex behaviour of BBHs in galactic nuclei influenced primarily by the tertiary SMBH, in the process adding to the TSUNAMI code the capability to produce strain signals. Our main conclusions are as follows:
@@
$\bullet$ In addition to the standard von Zeipel-Lidov-Kozai (ZLK) evolution where the inner argument of pericentre $\omega$ is circulating (Figure 3), $\omega$ and the eccentricity $e_{1}$ can also librate with large (Figure 4) or small (Figure 5) amplitude around a fixed point in phase space, which is given by Eq. (4) and $\omega=\pi/2,3\pi/2$ (see also Grishin, 2024a). Moreover, for extremely high eccentricity (or low $j_{z}$), GW emission at pericentre can lead to GW capture and merger even an initially librating orbit (Figure 6), which could potentially be observed both by LISA and ground-based GW detectors. These four families have unique signatures in $\omega$ and GW strain (Figure 2).
@@
$\bullet$ We added a method to calculate the strain polarisations within the N-body code TSUNAMI (Trani & Spera, 2023) by using the quadrupole formula and accounting for the direct accelerations from the tertiary and due to post-Newtonian effects up to 3.5PN. Using this method we are able to compare the different GW signatures directly from the simulations (Figure 8). We find that most of the strain signatures include high bursts separated by a quiescent phase on timescales of years. The small-amplitude orbit maintains its high eccentricity and has consistent burst on the order of its inner orbital period of a few days.
@@
$\bullet$ Due to the complexity of the dynamics, each orbit has almost unique signature. Varying the SMBH mass by as little as $3\%$ ($\gamma=0.99$ in Figure 11) will lead to an offset of the time-of-arrival of the burst by about $20$ s, which can still enable distinguishability, based on the precision to which we are able to isolate the arrival time of a single burst. Distinguishing BBH in galactic nuclei from those formed in more equal-mass field triples is also possible due to the breakdown of the test-particle limit and the backreaction on the tertiary orbit (Figure 11). Moreover, the orbits of BBHs with a SMBH perturber are markedly different from a binary in isolation with the same initial conditions (Figure 12).
$\bullet$ The GW signature should be distinguishable in the LISA band where the SNR is high enough (around 100 for peak frequency, Figure 9). Semi-analytical population studies suggest that $10\%$ of BBHs in the Galactic Centre (or 1000 BBHs in total) will have highly eccentric librating orbits (Figure 10).
@@
The duration of a burst is about $r_{p}/v_{p}$ at periapse, which is $\approx 1$ minute, as in Figure 11. This occurs every $\sim 3.4$ days, or >4800>4800 minutes. Generating 1 year of data takes $\approx 0.5$ seconds for output snapshots above $\Delta t=5,000$ sec, and linearly increases to $\sim 4$ minutes for the desired output of $\Delta t=10\ \rm s$ in order to capture the burst time derivative. Thus, full Bayesian parameter estimation is not feasible using the waveform model we have developed. However, searches and inference methods that rely on bursts and burst timing, rather than generation of full waveforms, can identify these sources and potentially infer the properties of the system (Romero-Shaw et al., 2023a; Knee et al., 2024; Saini et al., 2025).